OpenStudy (anonymous):

What are the solutions of the equation? c2-10c=0

jimthompson5910 (jim_thompson5910):

We can factor out 'c' to go from c^2 - 10c = 0 to c(c - 10) = 0 Now we can then use the zero product property to say c = 0 or c-10 = 0 Finally, solve that second equation for c to get c = 0 or c = 10